OpenStudy (anonymous):

-Under the Quadratic equation (ROOTS) Find the dimensions of a rectangle whose length is 5 less than twice its width and whose area is 63 square units.

OpenStudy (anonymous):

let width be w then the lenght = 2w - 5 area = l*w 63 = w(2w-5) 2w^2 - 5w - 63 = 0 2w^2 -14 w + 9w - 63 = 0 (2w+9)(w-7)=0 so width = 7 then length = 9